OpenStudy (anonymous):

When solving x^2 - 16x = 10 by completing the square what value is added to both sides of the equation? 6 8 16 64

OpenStudy (anonymous):

64

OpenStudy (anonymous):

x^2 -16x = 10 x^2 - 16x +(co-efficient of x/2)^2 =10 + (co-efficient of x/2)^2 Here , co-efficient of x is 16 So x^2 -16x + (16/2)^2 = 10 + (16/2)^2 x^2 -16x + (8)^2 = 10 + (8)^2 x^2 - 16 + 64 = 10 + 64 Therefore both sides add by 64